OpenStudy (anonymous):

How do I create a function with these characteristics? (pre-calc) Has rational coefficients and roots of 3, 3, and 3-i

OpenStudy (anonymous):

The 3-i is mainly what throws me off...

OpenStudy (jdoe0001):

well, the issue being... that say when you use the quadratic formula, and say you have a negative discriminant, then you'd end up with 2 COMPLEX solutions recall \(\bf \text{quadratic formula}\\ x= \cfrac{ - b \color{red}{\pm} \sqrt { b^2 -4ac}}{2a}\) what does that mean? well, it means that the complex root there, is not all by her lonesome, she's got a "conjugate" companion too, that is 3+i, so those are your roots

OpenStudy (jdoe0001):

\(\bf x=3\implies x-3=0\implies (x-3)=0\\ \quad \\ x=3\implies x-3=0\implies (x-3)=0\\ \quad \\ x=3-i\implies x-3+i=0\implies (x-3+i)=0\\ \quad \\ x=3+i\implies x-3-i=0\implies (x-3-i)=0\\ \quad \\ \quad \\ \color{green}{(x-3)(x-3)(x-3+i)(x-3-i)=0}\\ \quad \\ \quad \\ (x-3+i)(x-3-i)\implies [(x-3)-(i)][(x-3)+(i)]\\ \quad \\ \textit{recall that}\quad \color{blue}{(a-b)(a+b) = a^2-b^2}\qquad thus\\ \quad \\\ [(x-3)-(i)][(x-3)+(i)]\implies [(x-3)^2-i^2] \)

OpenStudy (jdoe0001):

and as you recall what \(\bf i^2\) is, then expand the binomial, cancel out like terms, and that'd be the binomial the complex came from

OpenStudy (jdoe0001):

the (x-3)(x-3) is just a multiple root, or a root with multiplicity of 2

OpenStudy (anonymous):

that all just confused me a lot /: do i just multiple (x - 3)(x - 3)(x - 3 - i)?

OpenStudy (anonymous):

*multiply

OpenStudy (jdoe0001):

hhehe

OpenStudy (jdoe0001):

ok... what part you didn't get?

OpenStudy (anonymous):

almost all of it lol

OpenStudy (jdoe0001):

have you even done quadratics at all? do you know how to factor them?

OpenStudy (anonymous):

I have, but still pretty new at them

OpenStudy (jdoe0001):

you know the difference of 2 squares, right? \(\bf a^2-b^2 = (a-b)(a+b)\)

OpenStudy (anonymous):

I recall something like that, yes

OpenStudy (jdoe0001):

... have you covered the quadratic formula?

OpenStudy (anonymous):

I have that written down in my notes

OpenStudy (anonymous):

so yes

OpenStudy (jdoe0001):

say for example, in a quadratic.... you end up with say using the quadratic formula and looks like \(\bf x= \cfrac{ - 5 \pm \sqrt { 3^2 -4(2)(3)}}{2(2)}\implies x= \cfrac{ - 5 \pm \sqrt { 9-24}}{4}\implies x= \cfrac{ - 5 \pm \sqrt {-15}}{4}\\ \quad \\ x=\cfrac{ - 5 + \sqrt {-15}}{4}\qquad and\qquad x=\cfrac{ - 5 - \sqrt {-15}}{4}\)

OpenStudy (jdoe0001):

as you can see, from there, you'd end up with 2, not 1, 2 imaginary or complex roots so the idea being that, above there you have 3 roots, but 1 is complex, but complex roots never come alone, they come with their conjugate companion so you really have 4 roots to work with, in order to get the function or polynomial

OpenStudy (jdoe0001):

hmm actually .... I .... 5 should have been 3... but anyhow, just to exemplify, that's the reason, why a complex will not come alone

OpenStudy (anonymous):

So then do I multiply (x-3)(x-3)(x-3-i)(x-3+i)?

OpenStudy (jdoe0001):

so the roots you were given PLUS the conjugate of the given complex will yield \(\bf \bf x=3\implies x-3=0\implies (x-3)=0\\ \quad \\ x=3\implies x-3=0\implies (x-3)=0\\ \quad \\ x=3-i\implies x-3+i=0\implies (x-3+i)=0\\ \quad \\ x=3+i\implies x-3-i=0\implies (x-3-i)=0\)

OpenStudy (jdoe0001):

yeap, which is what I showed above \(\bf \color{green}{(x-3)(x-3)(x-3+i)(x-3-i)=0}\\ \quad \\ \quad \\ (x-3+i)(x-3-i)\implies [(x-3)-(i)][(x-3)+(i)]\\ \quad \\ \textit{recall that}\quad \color{blue}{(a-b)(a+b) = a^2-b^2}\qquad thus\\ \quad \\\ [(x-3)-(i)][(x-3)+(i)]\implies [(x-3)^2-i^2]\)

OpenStudy (anonymous):

Ugh so many numbers lol it's overwhelming. So wait does that mean the equation would be [(x-3)^2 - i^2]?

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

Uggggh I'm sorrry lol

OpenStudy (anonymous):

I'm not usually this bad at math

OpenStudy (anonymous):

I'm giving you a medal cuz I know you sincerely are trying lol

OpenStudy (jdoe0001):

yes, lemme show it a bit more.... \(\bf \color{green}{(x-3)(x-3)(x-3+i)(x-3-i)=0}\\ \quad \\ \quad \\ (\color{red}{x-3}+\color{blue}{i})(\color{red}{x-3}-\color{blue}{i})\implies [(\color{red}{x-3})-(\color{blue}{i})][(\color{red}{x-3})+(\color{blue}{i})]\\ \quad \\ \textit{recall that}\quad \color{blue}{(a-b)(a+b) = a^2-b^2}\qquad thus\\ \quad \\\ [(x-3)-(i)][(x-3)+(i)]\implies [(\color{red}{x-3})^2-\color{blue}{i}^2]\)

OpenStudy (anonymous):

I just.... I don't understand what any of that means, like I can't comprehend it

OpenStudy (jdoe0001):

well.... do you know how I got \(\bf [(x-3)^2-i^2]\quad ?\)

OpenStudy (jdoe0001):

is really just plain terms grouping all I've done

OpenStudy (anonymous):

Yes, I think

OpenStudy (anonymous):

You used that rule

OpenStudy (anonymous):

But that's not the equation?

OpenStudy (jdoe0001):

the equation is obtained by multiplying its roots

OpenStudy (jdoe0001):

so far we've multiplied the 2 complex ones, so now we know that \(\bf (x-3)(x-3)(x-3+i)(x-3-i)=0\\ \quad \\ \implies (x-3)(x-3)[(x-3)^2-i^2]\)

OpenStudy (jdoe0001):

so... can you recall what \(\bf i^2\) is?

OpenStudy (anonymous):

Ohhh, I see! And i^2... uggggh I know it's 1 or -1 or something like that, right?

OpenStudy (jdoe0001):

\(\bf i^2\implies \sqrt{-1}\cdot \sqrt{-1}\implies (\sqrt{-1})^2\implies \sqrt{(-1)^2}\implies -1\)

OpenStudy (anonymous):

Okay, so it's -1

OpenStudy (jdoe0001):

\(\bf (x-3)(x-3)[(x-3)^2-i^2]\implies (x-3)(x-3)[(x-3)^2-(-1)]\\ \quad \\ (x-3)(x-3)[(x-3)^2+1]\implies (x-3)^2[(x-3)^2+1]\)

OpenStudy (jdoe0001):

so now all you have to do is, expand both binomials, and multiply their terms, then cancel any like-terms

OpenStudy (jdoe0001):

as you can see, it'd be a 4th degree polynomial, thus 4 degree, 4 roots

OpenStudy (anonymous):

well (x - 3)(x - 3) is x^2 - 6x + 9, and I could multiply that by itself (because there's another (x-3)^2) and then 1?

OpenStudy (jdoe0001):

well, you're right... but .... \(\bf (x-3)^2[(x-3)^2+1]\implies (x^2-6x+9)[x^2-6x+9+1]\\ \quad \\ (x^2-6x+9)[x^2-6x+10]\)

OpenStudy (anonymous):

I got x^4 - 12x^3 + 55x^2 - 114x + 90

OpenStudy (jdoe0001):

yeap, that's what I got :)

OpenStudy (anonymous):

yay! lol thank you so much for your time

OpenStudy (jdoe0001):

yw